Given a fixed amount of gas held at constant pressure, calculate the temperature to which the gas have to be changed if a 1.75 L sample at 23.0 degrees celsius were to have a final volume of 3.5 L


1
Expert's answer
2021-12-15T10:17:47-0500

"p=constant \\

V_1 = 1.75 \;L \\

T_1 = 23.0 + 273 = 296 \;K \\

V_2 = 3.5 \;L \\

\frac{V_1}{T_1} = \frac{V_2}{T_2} \\

T_2 = \frac{T_1V_2}{V_1} \\

T_2 = \frac{296 \times 3.5}{1.75} = 592 \;K \\

T_2 = 592 -273 = 319 \; °C"

Answer: 319 °C